• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name S. Sirca

    Log entry time 15:57:19 on March 5,2009

    Entry number 264757

    keyword=DAY shift summary


    DAY shift summary: S. Sirca (SL), H. Gao TO), D. McNulty (3rd)
    ==============================================================

    EVENTS
    ------
    08:00 took over at run #3978 from owl shift
    09:30 switch to kinematics G (HRSL down to 1.200 GeV/c)
    09:36 started new Compton (#16209, #60735) and parity (#30948) runs
    11:00 reference (N2) cell runs #2000 (BB) and #20547 (HRSL)
    11:40 back to production running with runs #2001/#20548
    11:59 take beam half-wave plate OUT
    12:40 a sequence of beam fluctuations trips all BB WC HVs
    13:51 beam OFF (arc magnet problems)

    RUNS
    ----

    Bigbite:

    07:25-08:19 #1994
    08:25-08:59 #1995
    09:01-09:29 #1996 (last run at kinematics I)
    09:33-10:17 #1997 (first run at kinematics G)
    10:20-10:58 #1998
    11:03-11:04 #1999 (junk run to set prescales)
    11:07-11:40 #2000 (N2 reference cell run)
    11:43-11:58 #2001
    11:59-12:40 #2002
    #2003 (not a production run: just to see the rates during beam-pos checks)
    13:48-13:51 #2004 (beam OFF)

    HRSL:

    07:25-09:30 #20545 (last run at kinematics I)
    09:38-11:01 #20546 (first run at kinematics G)
    11:03-11:40 #20547 (N2 reference cell run)
    11:43-11:58 #20548
    11:59-12:40 #20549
    #20550, 20551, 20552 (runs for spot++ checks)
    13:38-13:51 #20553 (beam off)

    Compton:

    until 09:36 #16208/#60374
    since 09:36 #16209/#60375
    since 11:32 #16210, since 11:41 #60376

    parity:

    until 09:36 #30947
    since 09:36 #30948
    since 11:42 #30949